karen took 5 h to drive 470 km from sudbury to brantford. for part of the trip, she drove at 100km/h. For the rest of the trip, she drove at 90 km/h. How far did she drive at each speed?
Answer by checkley71(8403)

100x+90(5-x)=470
100x+450-90x=470
10x=470-450
10x=20
x=2 hours she drove @ 100km/h.
5-2=3 hours she drove @ 90 km/h.
proof
100*2+90*3=470
200+270=470
470=470